MILLIMETER-WAVE SPECTROSCOPY AND COUPLED CLUSTER CALCULATIONS FOR NCCP

L. BIZZOCCHI, C. D. ESPOSTI, Dipartimento di Chimica “D. Ciamician”, Università di Bologna, Via F. Selmi 2, I-40126 Bologna, Italy; \underline{P. BOTSCHWINA} and B. SCHULZ, Institut für Physikalische Chemie, Universität Göttingen, Tammannstrasse 6, D-37077 Göttingen, Germany.